OpenStudy (anonymous):

Verify the identity: Sin(u+y+c)=sinucosycosc+cosusinycosc+cosucosysinc-sinusinysinc

OpenStudy (anonymous):

use sin(A+B)=sinA cosB+sinB cosA

OpenStudy (anonymous):

in your problem let A=u B=y+c

OpenStudy (zzr0ck3r):

\(\sin(a+b) = \sin(a)\cos(b)+\sin(b)\cos(a)\) \(\cos(a+b) = \cos(a)\cos(b)+\sin(a)\sin(b)\) so you have \(\sin(u+y+c) = \sin((u+y)+c) = \sin(u+y)\cos(c) + \sin(c)\cos(u+y) =\\ \cos(c)(\sin(u)\cos(y)+\sin(y)\cos(u))+\sin(c)(\cos(u)\cos(y)+\sin(u)\sin(y)) =\\ \small \cos(c)\sin(u)\cos(y)+\cos(c)\sin(y)\cos(u)+\sin(c)\cos(u)\cos(y)+\sin(c)\sin(u)\sin(y)\) as desired

OpenStudy (zzr0ck3r):

that was not fun

OpenStudy (anonymous):

Thank you so much. To both of you.